OpenStudy (anonymous):

integrate 4xsin^-1x dx

OpenStudy (psymon):

Looks like this would require integration by parts.

OpenStudy (luigi0210):

Btw @clairewee Welcome to Openstudy :)

OpenStudy (anonymous):

se integration by parts

OpenStudy (anonymous):

alternately use sin^-1x=z or x=sinz dx= cosz dz ur question becomes integrate 4zsinzcosz dz =integrate 2z*2sinzcosz dz=integrate 2zsin2zdz and then using integration by parts can be better

OpenStudy (anonymous):

@Luigi0210 thanks! i tried integrating by parts and got 2x^2sin^-1x - (integration of) 2x^2 /sqrt(1-x^2)

OpenStudy (anonymous):

@matricked so i change it to another form first before integrating by parts?

OpenStudy (anonymous):

2x^2 /sqrt(1-x^2) can be written as 2(1 - (1-x^2))/sqrt(1-x^2) =2*(1/sqrt(1-x^2) - sqrt(1-x^2)) now u have integrals for both .....

OpenStudy (anonymous):

its up to you

OpenStudy (anonymous):

@matricked thanks!

OpenStudy (anonymous):

yw hope both ways are clea to you